# 3VL: Using Trees to Improve Vision-Language Models’ Interpretability

### v1

@@ Abstract
Vision-Language models (VLMs) have proved effective at aligning image and text representations, producing superior zero-shot results when transferred to many downstream tasks. However, these representations suffer some key shortcomings in Compositional Language Concepts (CLC) understanding such as recognizing objects’ attributes, states, and relations between different objects. Moreover, VLMs typically have poor interpretability, making it challenging to debug and mitigate compositional-understanding failures. In this work, we introduce the Tree-augmented Vision-Language (3VL) model architecture and training technique accompanied by our proposed Anchor inference method and Differential Relevance (DiRe) interpretability tool. By expanding the text of an arbitrary image-text pair into a hierarchical tree structure using language analysis tools, 3VL allows inducing this structure into the visual representation learned by the model, enhancing its interpretability and compositional reasoning. Additionally, we show how Anchor, a simple technique for text unification, can be employed to filter nuisance factors while increasing CLC understanding performance, e.g., on the fundamental VL-Checklist benchmark. We also exhibit how DiRe, which performs a differential comparison between VLM relevancy maps, enables us to generate compelling visualizations of the reasons for a model’s success or failure.
@@ I Introduction
We propose a novel approach to address VLMs’ limitations in compositional understanding and interpretability by leveraging natural language hierarchical structures. Our approach is based on the idea of expanding the text of an image-text pair into a hierarchical tree structure using language analysis tools, and inducing this structure into the visual representation learned by the model. Specifically, we propose the Tree-augmented Vision-Language (3VL) model architecture and training technique, which allows for rich exploration of the text space using several levels of incremental text augmentation from coarse to fine-grained. Each level of the tree represents a progressive refinement of the text structure and captures an increasingly detailed aspect of the image-text relationship. We use this hierarchical structure to guide the learning of visual features in a way that improves the model’s compositional reasoning ability.
@@
We evaluate our approach on popular Compositional Language Concepts benchmarks and demonstrate that our method achieves state-of-the-art performance, outperforming existing VLM-based methods on these tasks. Furthermore, we show that our approach is not only effective but also interpretable, enabling us to both generate compelling visualizations of the reasons for model success or failure and effectively filter the image signal for increased performance.
@@ II Background
This section begins with a review of existing methods for interpretability of deep neural networks and then discusses the usage of trees for deep learning and how our approach differs from previous works. Finally, it describes existing efforts for compositional understanding in VLMs.
@@ II-A Interpretability of deep neural networks
Many approaches have been developed for explaining the decisions of neural networks [23, 21]. Most works interpret decisions of an already trained network. Gradient Saliency [6], GradCAM [7] and Integrated Gradients [9] use convolutional neural network gradients to find significant pixels in the input that impact the network output. LIME [8] and SHAP [10] segment first the input into super-pixels in order to find regions that explain the output. Wavelets [11] and Shearlets [12] have been used in order to further improve the region selection. In order to provide similar interpretability tools for transformers, Chefer et al. [24] proposed to use the concept of layer-wise Relevance Propagation [25], with gradient integration for the self-attention layers of transformers. HilaCAM [22] proposed a more general approach that can visualize transformer architectures more easily.
Another line of work focuses on ‘explainability by design’ pointing to some flaws [13, 14, 26, 16, 17, 18, 19, 20] in existing solutions for neural networks that are already trained. Instead, it is suggested to incorporate interpretable features in the model’s structure. One approach generates a set of linear transformations that are applied to the input [27]. Another strategy encourages the network to make decisions that are locally linear [28]. This linearity makes the decision more interpretable. Another method [29] generates a sequence of queries and makes a decision based on them. In [30] a variational model is used to optimize the queries usage. An alternative approach is to learn interpretable semantic concepts from the data and then make a decision based on them [31, 32, 33, 34, 35]. A similar technique teaches the network to output also a ‘why’ concept in addition to its decision [36]. Another work has incorporated a star model into network learning to get a relationship between locations in the query image and the decision of the network [37].
Figure 2: The tree loss and contrastive loss that are used for training 3VL. For the tree loss we first generate a caption tree and then sum the cross entropy loss in all tree levels. For the contrastive loss we calculate the average cross entropy loss over all image-text pairs in the batch.
@@ II-B Tree usage for deep learning
Trees have been used with neural networks in different constellations [38]. They have been incorporated into LSTMs [39] for improving semantic representations. Tree LSTMs have been combined with dependency parsing trees (DPTs) to improve visual grounding [40]. In [41, 42] decision trees were combined with the representation learning functionality of neural networks. In [43] tree were used with VLMs for improving their grammar induction in an unsupervised manner.
For improving the explainability of the network, a random forest has been trained on the network data and its decisions were used to regularize the neural network such that it generates axis aligned decisions [44]. Random forests have also been used in conjunction with pre-trained neural networks for improving the robustness of the latter to adversarial attacks [45, 46]. Another strategy that uses trees to improve interpretability is NBDT [47] that replaces the neural network final layer with a decision tree, which helps to understand better the neural network mistakes.
@@ II-C Compositionality in Vision-Language Models
One of the key challenges in developing effective vision and language models (VLMs) is achieving a compositional understanding of the underlying visual and linguistic elements.
@@
Despite these advances, recent studies like VL-CheckList [63], the Winoground Challenge [64], VSR [65] and COLA [66] have shown that VLMs still struggle with understanding fine-grained language details and Compositional Language Concepts (CLC). Another work has shown that there is a fundamental challenge for transformers with compositionality [67]. Therefore, there is still much work to be done in developing VLMs with stronger compositionality and robustness to complex, real-world scenarios.
Achieving a full understanding of the semantics of rich visual scenes involves the ability to detect individual entities, reason about their interactions and attributes, and ultimately understand the visual concepts within the scene. Structured representations have played a vital role in this process, and have been applied to a wide range of computer vision applications, such as vision and language [68, 69, 52, 70], scene graphs [71, 72, 73, 74, 75], relational reasoning [76, 77], human-object interactions [78, 79, 80], action recognition [81, 82, 83, 84, 85, 86, 87], and even image and video generation from graphs [88, 89, 90]. However, most of these works rely on detailed, manually curated supervision that often involves the annotation of location information and structural details. This results in limited-size or synthetic data sources for training, which can limit the effectiveness of the model. In [1], the authors have proposed a method for teaching CLC to large VLMs using only the available large-scale noisy Vision-Language data sources collected from the web, without the use of expensive manual curation. This approach aims to improve the scalability and effectiveness of VLMs’ understanding of Compositional Language Concepts.
Perhaps most similar to our augmentation method is the work of [1], which generates one negative caption per example by replacing one random word out of all possible candidates. Our proposed approach differs from the above by generating several augmentations per caption in the form of hierarchical trees for VLM training. Note that unlike previous works on compitional reasoning in VLMs, our approach aims at improving both compositional reasoning and interpretability.
@@ III-A Caption tree generation
  • At the third level we generate, for the positive text “several people standing in a green field”, the negative texts “several people gathered in a green field” and “several people standing out a green field”.
@@ III-B Tree-based training
CLIP contrastive loss. Given a batch of $N$ image-caption pairs we extract the image and text representations, $I_{r}$ and $T_{r}$, using CLIP’s image and text encoders. Then, we compute the pairwise cosine similarity scores $S_{j,k}$ for each image $j$ and caption $k$ in the batch. From the similarities matrix we calculate the cross entropy loss with softmax over the rows ($\mathcal{L}_{img}$) and the cross entropy loss with softmax over the columns ($\mathcal{L}_{txt}$). The final CLIP contrastive loss is the average of these two losses:
@@
Tree-based loss. For each image-caption pair, we first create a caption tree. Then, for each level of the tree we calculate the cosine similarity scores between the image and all captions at that level and calculate the Cross Entropy Loss. The final tree loss, $\mathcal{L}_{tree}$, is the sum of losses over all tree levels (Fig. 2).
@@
Training Details. We finetune OpenAI CLIP [48] ViT-B/32 [97] with rank=1 LoRA adapters on the training set of MS-COCO [54] for 12 epochs. We use AdamW optimizer with learning rate $63e-$ 6 and weight decay $0.1$ and train with a batch size of 64 on a single GeForce RTX 2080 Ti NVIDIA GPU. We performed a hyperparameter sweep and chose final parameters and number of epochs based on MS-COCO validation set.
@@
TABLE II: Top-1 accuracy on VL-Checklist VG Object dataset, with openAI CLIP VIT-B/32, NegCLIP [98], RB+LLM Negs [1], and 3VL. As an ablation we also compare to CLIP and CLIP+LoRA that were trained without tree-based loss.
@@
IV Relevancy Maps based Token Removal and Interpretability
IV-A Token Removal
Token Removal removes the least significant image tokens according to a given relevancy map. Then, the image with the removed tokens is used as input to the image encoder. We call this approach HilaCAM with Token Removal.
Using the relevancy maps generated for each image-text pair and Token Removal we point the model towards the more important parts of the image. By taking into account the relevancy maps of an image paired with a positive text and the same image paired with a negative text simultaneously, we are able to understand models’ decisions better. Combined with 3VL negatives tree generation we gain valuable insights into the underlying failure modes.
Note that removing less relevant tokens may enhance already existing biases. While Table VIII shows that this does not amplify the bias, further reduction of the model bias is still required.
TABLE III: Top-1 accuracy on VSR [65] benchmark, with openAI CLIP VIT-B/32, NegCLIP [98], RB+LLM Negs [1], and 3VL.
@@ IV-A Token Removal
IV-B HilaCAM Anchor
We generate image-text pairs relevancy maps using HilaCAM [22]. In our case where a single image has more than one possible text, it is natural to calculate a different relevancy map for each image-text pair. Each of these relevancy maps can be used for Token Removal. We refer from now on to this natural way of calculating multiple relevancy maps per text as HilaCAM as well.
Another way to use HilaCAM in cases with a single image and few texts is to form a new single text from the given input texts and use this new text to generate a single relevancy map. One way to form such a new single text is to use only parts of texts that differ from one another.
Having this Anchor text, we can generate a relevancy map from it and use it for Token Removal. Anchor with Token Removal is meant to focus both positive and negative texts on the same parts of the image. These parts of the image should contain the most important features for both positive and negative texts. Forcing a sort of cross-attention between the positive and negative texts.
IV-C Differential Relevance (DiRe)
Another way to get a single relevancy map when we have two input texts, is to generate one relevancy map per text and then generate a new relevancy map by subtracting the negative relevancy map from the positive one. Although this method is not “fair” as we use the knowledge of which caption is positive and which is negative, we can still leverage this method combined with Token Removal for interpretability. The reason for that is that DiRe provides importance scores to the image tokens and it does not ’directly intervene’ in the actual decision of the VLM. Therefore, if we find that the tokens of high relevancy according to DiRe correlate with better accuracy of the VLM , then we can get a better understanding of which tokens (that correspond to locations at the input image) affect the decision of the VLM and thus attain improved interpretability. In Figure 3 we show that indeed tokens of high relevancy according to DiRe correlate with better accuracy of the VLM (see details in Section V-D).
TABLE IV: COCO Image-Text retrieval with CLIP VIT-B/32, and 3VL
@@ IV-C Differential Relevance (DiRe)
TABLE V: FLICKR Image-Text retrieval with CLIP VIT-B/32 and 3VL
@@
TABLE VI: CLIPSeg segmentation with CLIP VIT-B/32, and 3VL
@@
TABLE VII: Number of failures on COCO testset per part of speech tag for vanilla CLIP, NegCLIP [98] and 3VL. 3VL improves by more than 50% on verbs(VERB) and by 46% on adpositions(ADP).
Model NOUN ADP VERB ADJ CLIP 3104 2927 1612 913 NegCLIP 2743 1989 900 695 3VL 1976 1575 780 612
TABLE VIII: CLIP fail rate on some pairs of positive-negative words on COCO testset.
Figure 4: A screen shot of some questions from the user study. The user-study instructions appear at the top left.

### v2

@@ Abstract
Vision-Language models (VLMs) have proven to be effective at aligning image and text representations, producing superior zero-shot results when transferred to many downstream tasks. However, these representations suffer from some key shortcomings in understanding Compositional Language Concepts (CLC), such as recognizing objects’ attributes, states, and relations between different objects. Moreover, VLMs typically have poor interpretability, making it challenging to debug and mitigate compositional-understanding failures. In this work, we introduce the architecture and training technique of Tree-augmented Vision-Language (3VL) model accompanied by our proposed Anchor inference method and Differential Relevance (DiRe) interpretability tool. By expanding the text of an arbitrary image-text pair into a hierarchical tree structure using language analysis tools, 3VL allows the induction of this structure into the visual representation learned by the model, enhancing its interpretability and compositional reasoning. Additionally, we show how Anchor, a simple technique for text unification, can be used to filter nuisance factors while increasing CLC understanding performance, e.g., on the fundamental VL-Checklist benchmark. We also show how DiRe, which performs a differential comparison between VLM relevancy maps, enables us to generate compelling visualizations of the reasons for a model’s success or failure. Our code is available at: https://github.com/niryellinek/3VL.
@@ I Introduction
We propose a novel approach to address VLMs’ limitations in compositional understanding and interpretability by leveraging natural language hierarchical structures. Our approach is based on the idea of expanding the text of an image-text pair into a hierarchical tree structure using language analysis tools and inducing this structure into the visual representation learned by the model. Specifically, we propose the Tree-augmented Vision-Language (3VL) model architecture and training technique, which allows for a rich exploration of the text space using several levels of incremental text augmentation from coarse to fine-grained. Each level of the tree represents a progressive refinement of the text structure and captures an increasingly detailed aspect of the image-text relationship. We use this hierarchical structure to guide the learning of visual features in a way that improves the model’s compositional reasoning ability.
@@
We evaluate our approach on popular Compositional Language Concepts benchmarks and demonstrate that our method achieves state-of-the-art performance, outperforming existing VLM-based methods on these tasks. Furthermore, we show that our approach is not only effective but also interpretable, enabling us to generate compelling visualizations of the reasons for model success or failure and effectively filter the image signal for increased performance.
By explainable in the paper we mean that our model emphasizes better the more important parts in the image, where the importance is correlated to human judgment. Thus, to assess that our model is more explainable according to this definition, we use token removal that checks the importance of patches remaining in the image and a user study that checks the correlation between what is important to humans and what is important to the model.
@@ II Background
This section begins with a review of existing methods for the interpretability of deep neural networks and then discusses the usage of trees for deep learning and how our approach differs from previous works. Finally, it describes existing efforts for compositional understanding in VLMs.
@@ II-A Interpretability of deep neural networks
Many approaches have been developed to explain the decisions of neural networks [23, 21]. Most works interpret decisions of an already trained network. Gradient Saliency [6], GradCAM [7], and Integrated Gradients [9] use convolutional neural network gradients to find significant pixels in the input that affect the network output. LIME [8] and SHAP [10] segment first the input into superpixels to find regions that explain the output. The work in [24] use large language models (LLMs) to interpret VLMs. Wavelets [11] and shearlets [12] have been used to further improve region selection. To provide similar interpretability tools for transformers, Chefer et al. [25] proposed to use the concept of layer-wise Relevance Propagation [26], with gradient integration for the self-attention layers of transformers. HilaCAM [22] proposed a more general method to visualize transformer architectures more easily. This method generates a relevancy map for each interaction between input modalities, whether through self-attention or bi-modal attention. Each relevancy map is derived from a forward pass through the attention layers and is averaged across the attention heads of each map using gradients.
Another line of work focuses on ‘explainability by design’ pointing to some flaws [13, 14, 15, 16, 17, 18, 19, 20] in existing solutions for neural networks that are already trained. For example, [13] shows that some saliency methods are independent of the model parameters and training data and thus cannot explain the failures of the models. Instead, it is suggested that interpretable features are incorporated into the structure of the model. One approach generates a set of linear transformations that are applied to the input [27]. Another strategy encourages the network to make decisions that are locally linear [28]. This linearity makes the decision more interpretable. Another method [29] generates a sequence of queries and makes a decision based on them. In [30] a variational model is used to optimize the queries usage. An alternative approach is to learn interpretable semantic concepts from the data and then make a decision based on them [31, 32, 33, 34, 35]. A similar technique teaches the network to output also a ‘why’ concept in addition to its decision [36]. StarNet [37] incorporates a star model into network learning to get a relationship between locations in the query image and the decision of the network. The back projection generates decision evidence heatmaps, effectively explaining the model’s decision. Our work differs from the above by incorporating both compelling visualizations of model decisions using HilaCAM and explainability by design using a fine-grained captions tree.
Figure 2: The tree loss and contrastive loss that are used for training 3VL. For the tree loss we first generate a caption tree and then sum the cross entropy loss in all tree levels. For the contrastive loss, we calculate the average cross-entropy loss over all image-text pairs in the batch.
@@ II-B Tree usage for deep learning
Trees have been used with neural networks in different constellations [38]. They have been incorporated into LSTMs [39] for improving semantic representations. Tree LSTMs have been combined with dependency parsing trees (DPTs) to improve visual grounding [40]. In [41, 42] decision trees were combined with the representation learning functionality of neural networks. In [43] trees were used with VLMs for improving their grammar induction in an unsupervised manner.
To improve the network’s explainability, a random forest has been trained on the network data. Its decisions were used to regularize the neural network such that it generates axis-aligned decisions [44]. Random forests have also been used in conjunction with pre-trained neural networks to improve the latter’s robustness to adversarial attacks [45, 46]. Another strategy that uses trees to improve interpretability is NBDT [47] which replaces the neural network final layer with a decision tree, which helps to understand better the neural network mistakes. In [48] the hierarchical representations of VLMs are analyzed and a method has been proposed to improve them.
Our work uses a tree both for achieving explainability by design and generating several levels of incremental text augmentations. The tree decomposes the sentences in a structured manner, which intuitively should help the network better understand the input structures and therefore make them more interpretable and easier to decompose. Moreover, our structuring improves the representations and grounding also with our negative augmentation method, which is much richer than previous works of negative augmentation. Note also that our tree structure helps inspecting VLM’s failures, e.g., as we show in Section V-F.
@@ II-C Compositionality in Vision-Language Models
One of the key challenges in developing effective VLMs is achieving a compositional understanding of the underlying visual and linguistic elements.
@@
Despite these advances, recent studies like VL-CheckList [64], the Winoground Challenge [65], VSR [66], VALSE [67] and COLA [68] have shown that VLMs still struggle with understanding fine-grained language details and CLC. In VL-Checklist [64] we have a set of images and a pair of positive and negative captions for each image. The model is tested for image-text retrieval. This benchmark aims to evaluate VLMs’ compositional understanding of objects, attributes, and relations. Each negative caption replaces one word for Objects, Relations, or Attributes in the positive sentence. Objects are tested for size and location invariance, the Attributes category checks for recognition of five different attributes: color, material, size, action, and state, and the Relations category checks for recognition of two relations between objects: action relation and spatial relation. Another work has shown that there is a fundamental challenge for transformers with compositionality [69]. Therefore, there is still much work to be done in developing VLMs with stronger compositionality and robustness to complex, real-world scenarios.
Achieving a full understanding of the semantics of rich visual scenes involves the ability to detect individual entities, reason about their interactions and attributes, and ultimately understand the visual concepts within the scene. Structured representations have played a vital role in this process, and have been applied to a wide range of computer vision applications, such as vision and language [52, 70, 53, 51], scene graphs [71, 72, 73, 74, 75], relational reasoning [76, 77], human-object interactions [78, 79, 80], action recognition [81, 82, 83, 84, 85, 86, 87], and even image and video generation from graphs [88, 89, 90]. Some works [91, 92, 93] take advantage of the compositional nature of natural language for solving VQA. However, most of these works rely on detailed, manually curated supervision that often involves the annotation of location information and structural details. This results in limited-size or synthetic data sources for training, which can limit the effectiveness of the model. In [1], the authors have proposed a method for teaching CLC to large VLMs using only the available large-scale noisy Vision-Language data sources collected from the web, without the use of expensive manual curation. This approach aims to improve the scalability and effectiveness of VLMs’ understanding of Compositional Language Concepts.
Perhaps most similar to our augmentation method is the work of [1], which generates one negative caption per example by replacing one random word out of all possible candidates. In [1] the authors prepare in advance a closed set of words for each VL-Checklist category. The possible candidates for replacement are all words that match one of the categories’ set of words. The negative word is chosen randomly from the corresponding set of words excluding the positive word itself. Our proposed approach differs from the above by generating several augmentations per caption in the form of hierarchical trees for VLM training and the use of WordNet and a LLM for the generation of each negative sample. Note that unlike previous works on compositional reasoning in VLMs, our approach aims at improving both compositional reasoning and interpretability.
@@ III-A Caption tree generation
  • At the third level we generate, for the positive text “several people standing in a green field”, the negative texts “several people sitting in a green field” and “several people standing out a green field”.
@@ III-B Tree-based training
CLIP contrastive loss. Given a batch of $N$ image-caption pairs we extract the image and text representations, $I_{r}$ and $T_{r}$, using CLIP’s image and text encoders. Then, we compute the pairwise cosine similarity scores $S_{j,k}$ for each image $j$ and caption $k$ in the batch. From the similarities matrix, we calculate the cross entropy loss with softmax over the rows ($\mathcal{L}_{img}$) and the cross entropy loss with softmax over the columns ($\mathcal{L}_{txt}$). The final CLIP contrastive loss is the average of these two losses:
@@
Tree-based loss. For each image-caption pair, we first create a caption tree. Then, for each level of the tree, we calculate the cosine similarity scores between the image and all captions at that level and calculate the Cross Entropy Loss. The final tree loss, $\mathcal{L}_{tree}$, is the sum of losses over all tree levels (Fig. 2).
@@
Training Details. We finetune OpenAI CLIP [49] ViT-B/32 [100] with rank=1 LoRA adapters on the training set of MS-COCO [55] for 12 epochs. We use an AdamW optimizer with a learning rate of $63e-$ 6 and weight decay $0.1$ and train with a batch size of 64 on a single GeForce RTX 2080 Ti NVIDIA GPU. We performed a hyperparameter sweep and chose the final parameters and the number of epochs based on the MS-COCO validation set.
IV Relevancy Maps based Token Removal and Interpretability
IV-A Token Removal
Token Removal removes the least significant image tokens according to a given relevancy map. We use a $7\times 7$ relevancy map which is generated by HilaCAM and holds a significance score per patch in the image. Thus, the image is divided also into $7\times 7$ equal-size patches, where each patch is an image token. The image with the removed tokens is used as input to the image encoder. We call this approach HilaCAM with Token Removal.
Using the relevancy maps generated for each image-text pair and Token Removal we point the model towards the more important parts of the image. By taking into account the relevancy maps of an image paired with a positive text and the same image paired with a negative text simultaneously, we are able to understand models’ decisions better. Combined with 3VL negatives tree generation we gain valuable insights into the underlying failure modes. See Section V-F for more details.
Figure 3: Generating one relevancy heatmap using an “Anchor” text from two text possibilites. Note that unlike Figure 4, instead of having two heatmaps, here we have only one heatmap that is generated from the ”Anchor” text that we create from the positive and negative texts.
Figure 4: When we have two possible texts for one image, we may apply HilaCAM two times to get two different relevancy heatmaps.
IV-B HilaCAM Anchor
HilaCAM [22] generates a relevancy heatmap for a given image-text pair. In cases where two text possibilities exist (e.g., one positive and one negative) for a single image, it is possible to generate one relevancy heatmap for each text paired with the given image. Formally, given an Image I, a positive text P, and a negative text N we apply HilaCAM two times to get two different relevancy heatmaps as illustrated in Figure 4. This provides us with the following two heatmaps:
$positive\_heatmap=HilaCAM(I,P)$ (3)
$negative\_heatmap=HilaCAM(I,N)$ (4)
By abuse of notation, we refer from now on to this natural way of calculating a different relevancy map per text as HilaCAM as well. Each such relevancy map can be used for a process we call ‘Token Removal’ that is described hereafter.
Instead of using these two heatmaps, we propose another way to use HilaCAM in cases with a single image and two texts, e.g., one positive and one negative. In our approach, denoted as “Anchor”, we form a new single text from the two input texts and use this new text to generate a single relevancy map. In cases where the negative text was formed by replacing a few words in the positive text, it is possible to form a new single text by focusing on parts of texts that differ from one another. We call this new text as the “Anchor” text.
Formally, given an Image I, a positive text P, and a negative text N (that was generated by replacing a few words in P). We apply Anchor on P and N to get the new “Anchor” text Anc. Then we apply HilaCAM on Anc to get a single Anchor relevancy heatmap. Figure 3 illustrates the use of HilaCAM with Anchor. The following series of formulas describe how we get the anchor heatmap:
[…]
AP = minimal noun or verb phrase in P that contains DP
AN = minimal noun or verb phrase in N that contains DN
Anc = Anchor (P, N) = AP + ‘ or ‘ + AN
anchor_heatmap = HilaCAM(I, Anc)
Having this Anchor text, we can generate a relevancy map from it and use it for Token Removal. Anchor with Token Removal is meant to focus both positive and negative texts on the same parts of the image. These parts of the image should contain the most important features for both positive and negative texts. Directing the network to focus on the common information between the positive and negative texts.
IV-C Differential Relevance (DiRe)
Another way to get a single relevancy map when we have two input texts is to generate one relevancy map per text and then generate a new relevancy map by subtracting the negative relevancy map from the positive one. Although this method is not “fair” as we use the knowledge of which caption is positive and which is negative, we can still leverage this method combined with Token Removal for interpretability. The reason for that is that DiRe provides importance scores to the image tokens and it does not ‘directly intervene’ in the actual decision of the VLM. Therefore, if we find that the tokens of high relevancy according to DiRe correlate with better accuracy of the VLM, then we can get a better understanding of which tokens (that correspond to locations at the input image) affect the decision of the VLM and thus attain improved interpretability. In Figure 5 we show that indeed tokens of high relevancy according to DiRe correlate with better accuracy of the VLM (see details in Section V-D).
@@ IV-C Differential Relevance (DiRe)
TABLE II: Top-1 accuracy on VL-Checklist VG Object dataset, with openAI CLIP VIT-B/32, NegCLIP [101], RB+LLM Negs [1], and 3VL. As an ablation, we also compare to CLIP and CLIP+LoRA which were trained without tree-based loss.
@@
TABLE III: Top-1 accuracy on VL-Checklist(VLC) Relation(Rel) average, Attribute(Attr) average, Object(Obj) average, and average zero-shot(ZS) on ELEVATER 21 datasets. We test on openAI CLIP VIT-B/32, NegCLIP [101], RB+LLM Negs [1], and 3VL. As an ablation, we also compare to CLIP and CLIP+LoRA that were trained without tree-based loss.
TABLE IV: Top-1 accuracy on VSR [66] benchmark, with openAI CLIP VIT-B/32, NegCLIP [101], RB+LLM Negs [1], and 3VL.
@@
TABLE V: COCO Image-Text retrieval with CLIP VIT-B/32, and 3VL
@@
TABLE VI: FLICKR Image-Text retrieval with CLIP VIT-B/32 and 3VL
@@
TABLE VII: CLIPSeg segmentation with CLIP VIT-B/32, and 3VL